OpenStudy (anonymous):

The total value of the pennies and nickels was $14.50. Hala counted the coins and found there were 450 coins in all. How many of each type of coin did she have?

OpenStudy (anonymous):

Use the standard form of a line: Ax+By=C Let .01=value of the pennies Let .05=value of the nickels . (.01)(x)+ (.05)(450-x)=14.50 .01x+22.50-.05x=14.50 [Solve for x] .01x-.05x=-22.50+14.50 -.04x=-8 x=200 pennies 450-200=250 nickels . Check by plugging the values back into the original equation: Ax+By=C (.01)(200)+(.05)(250)=$14.50 14.50=14.50 [Checks out]